OpenStudy (meilendurcer):

Bernadette bought two pieces of ribbon. One piece is 6 1/4 feet long, and the other is 11 1/2 feet long. Find the total length of the ribbon. i got 17 1/6

OpenStudy (anonymous):

I like to add the whole numbers first. So 6 + 11 = 17. And then add the fractions after they have the same common denominator. 2/4 + 1/4 = 3/4. I end up with 17 and 3/4

OpenStudy (anonymous):

Yeah, I also get the same answer with @michaeln. The way she gives also correct.

OpenStudy (meilendurcer):

oh ok.. thank you